OpenStudy (anonymous):

I am having trouble evluating: 2x1/10^1+3x1/10^2-4x1/10^3

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

1/10^1 = .1 1/10^2 = .01 1/10^3 = .001

OpenStudy (anonymous):

Thanks, answer would be 0.226